A Research Director of a certain university wants to replicate the result of the study 10 years ago with a standard deviation of 0.14. He wants to estimate the population mean to within an error of 0.04 of its true value. Using 95% confidence level, what is the sample size that he needs?


1
Expert's answer
2022-05-11T09:43:35-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} =1.96."

"E=z_c\times\dfrac{\sigma}{\sqrt{n}}\le0.04"

"n\ge(\dfrac{1.96\cdot0.14}{0.04})^2"

"n\ge48"